the sum of two numbers is 16. the difference is 4. what are the two numbers ?




Answer by ewatrrr(24785)


 
Hi
the sum of two numbers is 16. the difference is 4. what are the two numbers
Let x and y represent
Question states***
x+y =16
x-y = 4
2x = 20
x = 10 and y = 6 (x-y = 4 OR y = x-4)